### v1

@@
xujl@ihep.ac.cn
@@
Miao He
hem@ihep.ac.cn
Institute of High Energy Physics, Beijing, China
Cédric Cerna
cedric.cerna@in2p3.fr
Yongbo Huang
huangyb@gxu.edu.cn
Guangxi University, Nanning, China
@@ III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
@@ 1 Introduction
The Jiangmen Underground Neutrino Observatory (JUNO) is a particle physics experiment currently under construction, aiming for neutrino studies and exotic searches [1, 2, 3]. The central detector [4] is made of a 20 kiloton liquid scintillator target surrounded by 17,612 20-inch [5] and 25,600 3-inch photomultiplier tubes (PMTs) [6], all of which are submerged in water at a depth of up to 43.5 m along with their frontend electronics.
@@
The system is divided into 200 readout units of 128 PMTs each. A diagram of one such unit is shown in Fig. 1. Each PMT is equipped with a high voltage (HV) divider and a coaxial cable that can be 5 m or 10 m in length and that transmits both the HV and the signal. The PMT, HV divider, and one end of the cable are soldered and waterproofed. The other ends of sixteen cables are connected to a multichannel connector, each consisting of a plug and receptacle pair. These connectors are linked to an underwater stainless steel box (UWB), housing the 128-channel frontend power supply and readout electronics. This paper provides details on the instrumentation of the 3-inch PMTs, encompassing the design and production of the HV divider, cable, and connector, as well as the waterproof potting and the acceptance tests. The production and characterization of 26,000 (including 400 spares) 3-inch PMTs was reported in Ref. [6]. The electronics design was outlined in Ref. [3, 14, 15].
@@ 2 HV Divider
The HV divider follows a positive HV scheme. The HV ratio for the 11 series of dynodes is optimized for collection efficiency and timing measurements [16, 17, 18]. The resistor ratio for the first three dynodes is 3:2:1, and the other resistors are equal in value to the third one, 15 M (Fig. 2 (a)). The total resistance of 210 M and a working HV generally less than 1,300 V result in a gain of approximately and a direct current (DC) less than 6.5 µA. In consideration of safety and reliability requirements, the HV divider components include wire resistors, as depicted in Fig. 2 (b), due to superior voltage and thermal tolerance. The main challenge lies in the dense packing of components on the small PCB, which is designed to be soldered directly to the PMT pins and is constrained to have a similar size as the 48 mm diameter of the PMT glass bulb neck. Considering the re-evaluated requirements of the 3-inch PMT system for JUNO, the reliability requirement for the HV divider was slightly relaxed compared to the 20-inch PMT HV divider [5], allowing for the use of surface-mounted components. The final design is illustrated in Fig. 2 (c), where resistors of type 1206 with a maximum of 200 V and a 0.25 W power rating are selected, along with ceramic DC capacitors. The failure rate of the HV divider, as determined from reliability of individual components, is less than 0.05% per year.
@@
The HV divider is designed to achieve a maximum working HV of 2,000 V using a single layer of standard FR-4 [19] PCB material with a thickness of 1.2 mm and a diameter of 47 mm, in compliance with the radioactivity requirements of JUNO [20]. The PCB design ensures a clearance exceeding 1.6 kV/mm. Testing was performed on a few random samples of the final products, where HV dividers were subjected to several weeks of operation up to 2,400 V with satisfactory performance, and a maximum of 3,000 V was applied for a short period without failure. The flasher rate of the HV divider was verified to be less than 1 mHz, as detailed in Ref. [21]. To mitigate potential risks of flasher occurrences during production and long-term operation, black polyurethane was used to cover the divider PCB after it has been soldered to the PMT pins, as will be discussed in Sec. 4. A temperature cycling test was successfully conducted on 20 samples for 10 cycles without any failures, subjected to temperatures ranging from 5∘C to 55∘C for three hours per cycle. The signal quality coupled with the 3-inch PMT was further assessed in Ref. [22, 23], with additional results to be discussed in the following sections.
@@
Viking [24] surface bounding resistors of type 1206 and Murata [25] ceramic capacitors were chosen for the design. In adherence to the JUNO divider PCB procedure [21], mass production was executed after a successful small batch production of 100 pieces, all of which passed testing. All PCBs were soldered at the factory following cleaning at the designated soldering points. Subsequently, a visual inspection of the soldering points was performed. After another round of cleaning, each PCB underwent a 30-second maximum HV test at 2,000V before being soldered to the PMT and the cable. In total, 26,500 divider boards were produced which includes 900 spares. Acceptance tests were conducted during integrating with PMTs, as discussed in Sec. 4.2.
@@ 3.1 Design
To reduce costs and facilitate the installation of the JUNO detector, a novel 16-channel waterproof connector and cable was designed in collaboration with AXON’ company [26]. The connector’s schematic design is shown in Fig. 3 and consists of two parts: a plug that is attached to 16 PMTs through frontend cables, and a receptacle that is mounted on the UWB. Waterproof sealing between the 16 cables is achieved through injection molding in the plug and the application of adhesive in the receptacle. A double-sealing design ensures tightness when the plug is connected to the receptacle.
@@
A custom-made RG178 coaxial cable with a 50 impedance was also developed. The outer layer of the cable is made of High-Density Polyethylene (HDPE) with a diameter of 2.1 mm, designed to prevent fluorescence and ensure long-term reliability in ultra-pure water environment. As an additional safeguard, a water-blocking powder, developed by AXON’ company, was incorporated into the grounding layer. If the cable jacket is compromised, water would penetrate this layer, activating this barrier and preventing further water ingress. A water pressure test demonstrated that after one hour under 5 bar pressure, water infiltration was restricted to only 5 cm.
During the commissioning phase of the JUNO detector, a few weeks after the start of water filling, a significant number of channels experienced HV breakdown. This issue was traced to air diffusing from the connectors toward the ultra-pure degassed water, resulting in a pressure drop within the connector assembly. To address this problem, a continuous flow of nitrogen gas was introduced into the water circulation system, effectively resolving the issue. The 3-inch PMT system was integrated into detector operation following the completion of liquid scintillator filling. The operating power of each PMT group is being continuously monitored and has remained stable. Further details and comprehensive analysis of this incident will be provided in an upcoming publication.
Figure 3: (a) Schematic design of the 16-channel connector consisting of a plug (left part) and a receptacle (right part). The red dots and lines indicate water-resistant O-rings or gaskets. Due to intellectual property rights, further details cannot be disclosed. (b) A picture of a connector set with cables.
3.2 Signal Performances
The quality of the PMT signal was examined using a prototype of the connector and cables, as detailed in Ref. [23]. Typical signals for the single photoelectron (SPE) with a gain of 3106 are displayed in Fig. 4 (a). Small reflections following the main signal, attributed to a slight impedance mismatch not to the connector but in the hand-made HV-signal splitter, were observed. This issue was resolved in the final version of the splitter as reported in Ref. [15]. To compensate for the loss of PMT gain due to cable attenuation, an additional 23 V or 45 V voltage was applied for 5 m or 10 m cables, respectively. The observed amplitude, rise time, fall time, and Full Width at Half Maximum are listed in Tab. 1.
@@ 3.3 Long Term Validation
The JUNO 3-inch PMT system was designed to achieve a channel failure rate lower than 10% over a 6-year period, equally distributed among the PMTs and their readout electronics. This translates into an estimated Mean Time Between Failure (MTBF) exceeding 120 years, or equivalently, a Failure In Time (FIT) value below 1,000. This requirement applies directly to the cables and connectors that interface with the photomultipliers. Estimating such a low failure rate is challenging when dealing with a limited number of components over a short testing period. For instance, evaluating the reliability of two underwater connector plug and receptacle systems would necessitate a 40-year assessment.
In ultra-pure water, the primary cause of long-term failure is chemical degradation, which compromises the protective barriers against water ingress such as cable jackets and connector gaskets. The test duration can be accelerated by subjecting the system to increased temperatures in a controlled setting. Compared to JUNO’s operational water temperature of C, raising the temperature by an additional C can expedite the testing process by a factor of approximately 2.5. Thus, a test conducted at C for 12 months simulates the effect of 40 years of ageing at the normal operational temperature of C.
@@
During the two testing phases conducted in France, spanning over 11,000 hours, no leaking current was observed, and both the cable and connector functioned perfectly, achieving the required reliability with a FIT value smaller than 500. During the destructive test conducted in Chile, five receptacles were aged to the equivalent of over 120 years and exhibited no leaks, yielding a FIT value of 350.
@@ 3.4 Production and Quality Inspection in Axon’
Approximately 200 km of cables were produced and thoroughly tested, with 100% of the insulation verified through a 3,000 V spark test. After the cabling and overmolding processes were finalized, each plug was tested with 1.8 kV of HV for 70 seconds to confirm functionality. Subsequently, the plugs were equipped with stainless steel accessories that matched those of the receptacle. The sealing was achieved using a double sealing design.
The sealed cables and plugs were placed into a stainless steel tank filled with ultrapure water, as shown in Fig. 5. Humidity test paper was also added. The tank was closed, and the water pressure was raised to 0.75 MPa, maintained for two hours. Once removed from the water tank and dried, the water sealing was verified through visual inspection of the humidity test paper and a subsequent HV test. All cables and connectors underwent testing. Those passing the tests were then used to create the required 10 m and 5 m segments.
@@
For each batch, the first and last receptacles were mounted on the lid of the water tank with the 16 pins to be submerged in water (Fig. 5). If these two receptacles passed the test, the remaining receptacles from the batch proceeded to the next step, which involved attaching the micro coaxial connector to the other end of the cable that connects to the electronics in the UWB. In the case of failure, a more thorough inspection and additional destructive pressure testing were conducted on the batch in question. Finally, the receptacle underwent the standard HV test at 1.8 kV for 70 seconds.

### v2

@@
Design, waterproofing, and mass production of the 3-inch PMT frontend system of JUNO
@@
Corresponding authors. Email addresses: xujl@ihep.ac.cn (Jilei Xu), hem@ihep.ac.cn (Miao He), cedric.cerna@in2p3.fr (Cédric Cerna), huangyb@gxu.edu.cn (Yongbo Huang).
@@
Miao He∗Institute of High Energy Physics, Beijing, China
Cédric Cerna∗Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Yongbo Huang∗Guangxi University, Nanning, China
@@ III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@
Technische Universität München, München, Germany
@@ 1 Introduction
The Jiangmen Underground Neutrino Observatory (JUNO) is a particle physics experiment aiming for neutrino studies and exotic searches [11, 18, 4]. The experiment started collecting physics data on August 26, 2025, and the initial detector performance results were documented in Ref. [9], with the first report on reactor neutrino oscillations measurement published in Ref. [8]. The central detector [7] is made of a 20 kiloton liquid scintillator target surrounded by 17,612 20-inch [1] and 25,600 3-inch photomultiplier tubes (PMTs) [14], all of which are submerged in water at a depth of up to 43.5 m along with their frontend electronics.
@@
The system is divided into 200 readout units of 128 PMTs each. A diagram of one such unit is shown in Fig. 1. Each PMT is equipped with a high voltage (HV) divider and a coaxial cable that can be 5 m or 10 m in length and that transmits both the HV and the signal. The PMT, HV divider, and one end of the cable are soldered and waterproofed. The other ends of sixteen cables are connected to a multichannel connector, each consisting of a plug and receptacle pair. These connectors are linked to an underwater stainless steel box (UWB), housing the 128-channel frontend power supply and readout electronics. This paper provides details on the instrumentation of the 3-inch PMTs, encompassing the design and production of the HV divider, cable, and connector, as well as the waterproof potting and the acceptance tests. The production and characterization of 26,000 (including 400 spares) 3-inch PMTs were carried out at Hainan Zhanchuang Photonics Technology Co., Ltd (HZC) [27] and reported in Ref. [14]. The electronics design was outlined in Ref. [4, 17, 32].
@@ 2 HV Divider
The HV divider follows a positive HV scheme. The HV ratio for the series of 11 dynodes is optimized for collection efficiency and timing measurements [34, 33, 15]. The resistor ratio for the first three dynodes is 3:2:1, and the other resistors are equal in value to the third one, 15 M (Fig. 2 (a)). The total resistance of 210 M and a working HV generally less than 1,300 V result in a gain of approximately and a direct current (DC) less than 6.5 µA. In consideration of safety and reliability requirements, the HV divider components include wire resistors, as depicted in Fig. 2 (b), due to superior voltage and thermal tolerance. The main challenge lies in the dense packing of components on the small printed circuit board (PCB), which is designed to be soldered directly to the PMT pins and is constrained to have a similar size as the 48 mm diameter of the PMT glass bulb neck. Considering the re-evaluated requirements of the 3-inch PMT system for JUNO, the reliability requirement for the HV divider was slightly relaxed compared to the 20-inch PMT HV divider [1], allowing for the use of surface-mounted components. The final design is illustrated in Fig. 2 (c), where resistors of type 1206 with a maximum of 200 V and a 0.25 W power rating are selected, along with ceramic DC capacitors. The failure rate of the HV divider, as determined from reliability of individual components, is less than 0.05% per year.
@@
The HV divider is designed to achieve a maximum working HV of 2,000 V using a single layer of standard FR-4 [25] PCB material with a thickness of 1.2 mm and a diameter of 47 mm, in compliance with the radioactivity requirements of JUNO [3]. The PCB design guarantees a minimum clearance of 1.6 kV/mm between neighboring printed wires with varying voltages. Testing was performed on a few random samples of the final products, where HV dividers were subjected to several weeks of operation up to 2,400 V with satisfactory performance, and a maximum of 3,000 V was applied for a short period without failure. The flasher rate, which refers to the frequency of HV luminous microdischarges in the laboratory air of the HV divider, was verified to be less than 1 mHz, as detailed in Ref. [37]. To mitigate potential risks of flasher occurrences during production and long-term operation, black polyurethane was used to cover the divider PCB after it has been soldered to the PMT pins, as will be discussed in Sec. 4. A temperature cycling test was successfully conducted on 20 samples for 10 cycles without any failures, subjected to temperatures ranging from 5∘C to 55∘C for three hours per cycle. The signal quality coupled with the 3-inch PMT was further assessed in Ref. [29, 36], with additional results to be discussed in the following sections.
@@
Viking [26] surface mounting resistors of type 1206 and Murata [24] ceramic capacitors were chosen for the design. In adherence to the JUNO divider PCB procedure [37], mass production was executed after a successful small batch production of 100 pieces, all of which passed testing. All PCBs were soldered at the factory following cleaning at the designated soldering points. Subsequently, a visual inspection of the soldering points was performed. After another round of cleaning, each PCB underwent a 30-second maximum HV test at 2,000V before being soldered to the PMT and the cable. In total, 26,500 divider boards were produced including 900 spares. Acceptance tests were conducted during integration with PMTs, as discussed in Sec. 4.2.
@@ 3.1 Design
To reduce costs and facilitate the installation of the JUNO detector, a novel 16-channel waterproof connector and cable was designed in collaboration with AXON’ company [21]. The connector’s schematic design is shown in Fig. 3 and consists of two parts: a plug that is attached to 16 PMTs through frontend cables, and a receptacle that is mounted on the UWB. Waterproof sealing between the 16 cables is achieved through injection molding in the plug and the application of adhesive in the receptacle. A double-sealing mechanism relying on redundant axial and radial o-rings ensures tightness when the plug is connected to the receptacle.
Further design details and descriptions can be found in the Chinese patent [30] held by AXON’.
@@
A custom-made RG178 coaxial cable with a 50 impedance was also developed. The outer layer of the cable is made of High-Density Polyethylene (HDPE) with a diameter of 2.1 mm, designed to prevent fluorescence and ensure long-term reliability in the ultra-pure water environment. To provide an extra layer of protection, the grounding layer was enhanced with a water-blocking powder based on sodium polyacrylate, a superabsorbent polymer developed by AXON’.
If the cable jacket is compromised, water would penetrate this layer, activating this barrier and preventing further water ingress. A water pressure test demonstrated that after one hour under 5 bar pressure, water infiltration was restricted to only 5 cm.
During the commissioning phase of the JUNO detector, a few weeks after the start of water filling, a significant number of channels experienced HV breakdown. The similar phenomenon was also observed in the SNO detector two decades ago [12]. This issue was traced to air diffusing from the connectors toward the ultra-pure degassed water. This occurred because the air concentration within the connector exceeded that of the degassed water, leading to a drop in pressure within the connector assembly. Since the insulation between the core and the ground of coaxial pins is provided by the residual air within the connector at atmospheric pressure, the breakdown voltage decreases as the pressure decreases according to Paschen’s Law [20].
To address this problem, a continuous flow of nitrogen gas was introduced into the water circulation system, which increase the overall gas concentration to levels higher than inside the connector. As a result, nitrogen gas diffuses from the water into the connector. This elevated the overall gas concentration to levels surpassing those inside the connector. Consequently, nitrogen gas diffused from the water into the connector, raising the pressure inside.
This increase in pressure consequently raises the breakdown voltage, ensuring it exceeds the operational high voltage requirement. The 3-inch PMT system was integrated into the detector operation following the completion of liquid scintillator filling. The operating power of each PMT group is being continuously monitored and has remained stable. Further details and comprehensive analysis of this issue and its solution will be provided in an upcoming publication.
Figure 3: (a) Schematic design of the 16-channel connector consisting of a plug (left part) and a receptacle (right part). The red dots and lines indicate water-resistant O-rings or gaskets. Further details are protected by intellectual property rights [30]. (b) A picture of a connector set with cables.
3.2 Signal Performance
The quality of the PMT signal was examined using a prototype of the connector and cables, as detailed in Ref. [36]. Typical signals for the single photoelectron (SPE) with a gain of 3106 are displayed in Fig. 4 (a). Small reflections following the main signal, attributed to a slight impedance mismatch not to the connector but in the hand-made HV-signal splitter, were observed. This issue was resolved in the final version of the splitter as reported in Ref. [32]. To compensate for the loss of PMT gain due to cable attenuation, an additional 23 V or 45 V voltage was applied for 5 m or 10 m cables, respectively. The observed amplitude, rise time, fall time, and Full Width at Half Maximum (FWHM) are listed in Tab. 1.
@@ 3.3 Long Term Validation
The JUNO 3-inch PMT system was designed to achieve a channel failure rate below 10% over a 6-year period, equally allocated between the PMTs and their readout electronics. Under the assumption of a constant failure rate described by an exponential lifetime distribution, a 5% failure probability over six years for the electronics corresponds to a Mean Time Between Failures (MTBF) of 117 years. In practice, this requirement is implemented by assuming an MTBF exceeding 120 years, which is equivalent to a Failure In Time (FIT) value below 951. This requirement applies directly to the cables and connectors that interface with the photomultipliers. Estimating such a low failure rate is challenging when dealing with a limited number of components over a short testing period. For instance, evaluating the reliability of two underwater connector plug and receptacle systems would necessitate a 40-year assessment.
In ultra-pure water, the primary cause of long-term failure is chemical degradation, which compromises the protective barriers against water ingress such as cable jackets and connector gaskets. The test duration can be accelerated by subjecting the system to increased temperatures in a controlled setting. Compared to JUNO’s operational water temperature of C, raising the temperature by an additional C can expedite the testing process by a factor of approximately 2.5 [28]. Thus, a test conducted at C for 12 months simulates the effect of 40 years of ageing at the normal operational temperature of C.
@@
During the two testing phases conducted in France, spanning over 11,000 hours, no leaking current was observed, and both the cable and connector functioned perfectly, achieving the required reliability with a FIT value smaller than 500. During the destructive test conducted in Chile, five receptacles were aged to the equivalent of over 120 years and exhibited no leaks, yielding a FIT value of 300.
@@ 3.4 Production and Quality Inspection in Axon’
Approximately 200 km of cables were produced and thoroughly tested, with 100% of the insulation integrity verified through a 3,000 V spark test during the cable production process. Subsequently, the cable was cut to specific lengths, 10 m and 20 m for the plugs, and 25 cm for the receptacles.
A bundle of 16 cables, either 10 m or 20 m long, were equipped with plugs at both ends and exposed to 1.8 kV voltage for 70 seconds to validate their operation. Cables that successfully passed this evaluation were enclosed by two imitation receptacles and immersed in an ultrapure water-filled stainless steel tank, as shown in Fig. 5. Moisture test paper was included within the connector. The tank was closed, and the water pressure was raised to 0.75 MPa, sustained for two hours. Following removal from the water tank and drying, the effectiveness of water sealing was confirmed via a visual examination of the moisture test paper and a subsequent HV test. All cables and connectors underwent rigorous testing procedures. Those that met the specified criteria were then bisected at the midpoint to create the required 10 m and 5 m segments.
@@
For each batch, the first and last receptacles were mounted on the lid of the water tank with the 16 pins to be submerged in water at a pressure of 0.75 MPa and sustained for two hours (refer to Fig. 5).
This testing procedure was classified as destructive due to the exposure of the pins to water. If the two receptacles mentioned above successfully passed this test, the remaining receptacles within the same batch proceeded to the subsequent stage. This stage involved connecting a micro coaxial connector to the opposite end of the cable, which links to the electronics in the UWB. In the case of a failure, all other receptacles in the batch underwent a similar underwater test, but this time sealed with a plug. Subsequently, all receptacles were subjected to the standard HV test at 1.8 kV for a duration of 70 seconds.